# An LLM Benchmark for Addressee Recognition in Multi-modal Multi-party Dialogue

### v1

@@ 1 Introduction
Figure 1: Snapshot from TEIDAN corpus
@@
session-01-city 6:14 065 / 081 / 137 13 / 17 / 13 11 32
session-02-city 5:50 076 / 094 / 098 16 / 22 / 24 15 47
@@
session-04-city 5:46 146 / 142 / 123 33 / 22 / 28 11 71
session-05-city 5:18 108 / 119 / 095 45 / 44 / 48 37 100
All / Ave. 29:20 / 5:52 - - 80 / 16 322 / 64.4
@@
  • A new corpus of spontaneous, multi-modal, triadic dialogues, offering a unique resource for this understudied area
  • The first LLM benchmark specifically for addressee recognition in multi-modal, multi-party dialogue, highlighting the challenges and need for further innovation
Ultimately, this research aims to establish a strong foundation for the development of advanced multi-party dialogue systems capable of understanding and responding appropriately in complex, real-world conversational settings.
@@ 2 TEIDAN Corpus
We first briefly describe the TEIDAN multi-party corpus. The corpus is of free discussion, as opposed to other works where a specific setting was used such as a meeting Carletta (2007); Mostefa et al. (2007) or the participants were involved in a task Kontogiorgos et al. (2018); Nihei et al. (2014) or game Stefanov and Beskow (2016); Litman et al. (2016); Hung and Chittaranjan (2010). Other multiparty corpora also exist for online discussions Reverdy et al. (2022).
Discussions consisted of participants in a triad (three people). Each participant was seated in a circle with a table in the center, as shown in Figure 1. Cameras captured the face of each participant while separate pin microphones were attached to each of them.
We had three general topics of conversation: which city would be best for the alternative capital of Japan, which items would be necessary to bring to a desert island, and where they would like to travel on the weekend. Each triad conducted the discussions three times, corresponding to the above topics.
Triads conversed for approximately 5-10 minutes per session, with no requirement to reach a conclusion. Data was collected from 10 sets of triads, resulting in a total of 30 discussion sessions. Note that this corpus is in the Japanese language.
@@ 3 Annotation of Addressee
(1) Initially, turns and the current speaker were annotated. Since the original TEIDAN corpus contains only IPU utterance segments, turn segments within the dialogue were annotated by removing certain utterances, including backchannels. This ensured that only one speaker could hold the floor at any given time. Minimal overlap was permitted during turn transitions.
@@
We have so far annotated five sessions from the TEIDAN corpus. Table 1 summarizes the annotation results. The analysis revealed that only approximately 20% (80 / 322) of turns explicitly specify an addressee. The ‘O’ label, indicating no specific addressee, was prevalent, particularly in discussions involving multiple opinions (statements). This result implies that a multi-party dialogue system that participates in this type of discussion and disregards addressee information may potentially interrupt the dialogue in 20% of turn-taking instances, assuming that they can always correctly recognize the end of the turn of a human participant.
@@ 4 Benchmark for Addressee Recognition
The GPT-4o achieved an accuracy of 80.9%, which is only marginally above chance level (80.1%). This indicates that the model struggles to identify the addressee in multi-party dialogues. The output by the LLM is summarized in Table 2 which shows that the model tends to output ‘O’, indicating that it often fails to recognize when an utterance is directed at a specific participant.
@@ 6 Benchmark for Next Speaker Prediction
We then evaluated GPT-4o’s performance on this next speaker prediction task. Using a prompt similar to that used for addressee recognition, the model was tasked with predicting the actual next speaker. The output label was limited to A, B, or C, with a chance-level accuracy of 50%, as either of the other two participants could take the turn. As a result, GPT-4o attained an accuracy of 46.0% on this task, performing below the chance level. This outcome further suggests that the model struggles to effectively capture the dynamics of turn-taking in spontaneous multi-party dialogues.
@@ 7 Conclusion
This study investigated the challenges of addressee and next speaker prediction in multi-party dialogues. We introduced a new multi-party dialogue corpus and analyzed the performance of an LLM (GPT-4o) on these tasks. The findings revealed that LLMs struggle with the complexities of multi-party interactions. They perform only marginally above the chance level in addressee recognition and below the chance level in the next speaker prediction task. Although the LLM was given the simple gaze feature, it did not improve the performance.

### v2

@@ Abstract
 [^0] This paper has been accepted for presentation at International Workshop on Spoken Dialogue Systems Technology 2025 (IWSDS 2025) and represents the author’s version of the work.
@@ 1 Introduction
Figure 1: A snapshot from TEIDAN corpus
@@
session-01-city 6:14 065 / 081 / 137 12 / 16 / 13 9 32
session-02-city 5:50 076 / 094 / 098 16 / 22 / 23 14 47
@@
session-04-city 5:46 146 / 142 / 123 29 / 21 / 24 10 64
session-05-city 5:18 108 / 119 / 095 44 / 43 / 46 36 97
Total (Ave.) 29:20 (5:52) 1616 (323.2) 387 (77.4) 75 (15) 312 (62.4)
@@
  • The introduction of the TEIDAN corpus, a new dataset of spontaneous, multi-modal, triadic dialogues that provides a unique resource for this understudied area
  • The development of the first LLM benchmark specifically designed for addressee recognition in multi-modal, multi-party dialogue, underscoring the challenges and the need for further innovation
Ultimately, this research aims to establish a strong foundation for the development of advanced multi-party dialogue systems capable of understanding and responding effectively in complex, real-world conversational settings.
@@ 2 TEIDAN Corpus
We begin by briefly describing the TEIDAN multi-party corpus. Unlike other datasets that involve specific contexts, such as meetings Carletta (2007); Mostefa et al. (2007), task-oriented interactions Kontogiorgos et al. (2018); Nihei et al. (2014), or game-based scenarios Stefanov and Beskow (2016); Litman et al. (2016); Hung and Chittaranjan (2010), the TEIDAN corpus captures goal-free discussions. Additional multi-party corpora also exist for online discussions Reverdy et al. (2022).
The discussions involved triads (groups of three participants). Each participant was seated in a circle around a table placed at the center, as shown in Figure 1. Cameras recorded each participant’s face, and individual pin microphones were used to capture their speech separately.
Participants discussed three general topics: (1) which city would be best suited as an alternative capital of Japan, (2) which items would be essential to bring to a desert island, and (3) where they would like to travel on the weekend. Each triad completed three discussion sessions, one for each topic.
Each triad conversed for approximately 5 to 10 minutes per session, with no requirement to reach a conclusion. Data were collected from 10 triads, resulting in a total of 30 discussion sessions. Note that this corpus is in the Japanese language.
@@ 3 Annotation of Addressee
(1) Initially, turns and the current speaker were annotated. Since the original TEIDAN corpus contains only IPU (inter-pausal unit) utterance segments, turn segments within the dialogue were annotated by removing certain utterances, including backchannels. This ensured that only one speaker could hold the floor at any given time. Minimal overlap was permitted during turn transitions.
@@
We have so far annotated five sessions from the TEIDAN corpus. Table 1 summarizes the annotation results. The analysis revealed that only approximately 20% (75 / 387) of turns explicitly specify an addressee. The ‘O’ label, indicating no specific addressee, was prevalent, particularly in discussions involving multiple opinions (statements). This result implies that a multi-party dialogue system that participates in this type of discussion and disregards addressee information may potentially interrupt the dialogue in 20% of turn-taking instances, assuming that they can always correctly recognize the end of the turn of a human participant.
@@ 4 Benchmark for Addressee Recognition
The GPT-4o achieved an accuracy of 80.9%, which is only marginally above the chance level (80.6%). This indicates that the model struggles to identify the addressee in multi-party dialogues. The output by the LLM is summarized in Table 2 which shows that the model tends to output ‘O’, indicating that it often fails to recognize when an utterance is directed at a specific participant.
@@ 6 Benchmark for Next Speaker Prediction
We then evaluated GPT-4o’s performance on this next speaker prediction task. Using a prompt similar to that used for addressee recognition, the model was tasked with predicting the actual next speaker. The output label was limited to A, B, or C, with a chance-level accuracy of 50%, as either of the other two participants could take the turn. As a result, GPT-4o attained an accuracy of 46.0% on this task, performing below chance level. This outcome further suggests that the model struggles to effectively capture the dynamics of turn-taking in spontaneous multi-party dialogues.
@@ 7 Conclusion
This study investigated the challenges of addressee and next speaker prediction in multi-party dialogues. We introduced a new multi-party dialogue corpus and analyzed the performance of an LLM (GPT-4o) on these tasks. The findings revealed that LLMs struggle with the complexities of multi-party interactions. They perform only marginally above chance level in addressee recognition and below chance level in the next speaker prediction task. Although the LLM was given the simple gaze feature, it did not improve the performance.